# Quantitative evaluation of methods to analyze motion changes in single-particle experiments

### v2

@@ Methods
[…]
Data availability
Code availability
All software used for the Challenge is available at https://github.com/AnDiChallenge/andi_datasets.
Acknowledgements.
G.M-G. acknowledges support from the European Union (ERC, QuantAI, Project No. 10105529) and the Austrian Science Fund (FWF) through the SFB BeyondC F7102. M.L. acknowledges support from: ERC AdG NOQIA; Ministerio de Ciencia y Innovación Agencia Estatal de Investigaciones (PGC2018-097027-B-I00/10.13039/501100011033, CEX2019-000910-S/10.13039/501100011033, Plan Nacional FIDEUA PID2019-106901GB-I00, FPI, QUANTERA MAQS PCI2019-111828-2, QUANTERA DYNAMITE PCI2022-132919, Proyectos de I+D+I “Retos Colaboración” QUSPIN RTC2019-007196-7); MICIIN with funding from European Union NextGenerationEU(PRTR-C17.I1) and by Generalitat de Catalunya; Fundació Cellex; Fundació Mir-Puig; Generalitat de Catalunya (European Social Fund FEDER and CERCA program, AGAUR Grant No. 2021 SGR 01452, QuantumCAT \  U16-011424, co-funded by ERDF Operational Program of Catalonia 2014-2020); Barcelona Supercomputing Center MareNostrum (FI-2022-1-0042); EU Horizon 2020 FET-OPEN OPTOlogic (Grant No 899794); EU Horizon Europe Program (Grant Agreement 101080086 — NeQST), National Science Centre, Poland (Symfonia Grant No. 2016/20/W/ST4/00314); ICFO Internal “QuantumGaudi” project; European Union’s Horizon 2020 research and innovation program under the Marie-Skłodowska-Curie grant agreement No 101029393 (STREDCH) and No 847648 (“La Caixa” Junior Leaders fellowships ID100010434: LCF/BQ/PI19/11690013, LCF/BQ/PI20/11760031, LCF/BQ/PR20/11770012, LCF/BQ/PR21/11840013). Views and opinions expressed in this work are, however, those of the author(s) only and do not necessarily reflect those of the European Union, European Climate, Infrastructure and Environment Executive Agency (CINEA), nor any other granting authority. Neither the European Union nor any granting authority can be held responsible for them. R.M. acknowledges DFG grant ME 1535/12-1. D.K. acknowledges funding from the National Science Foundation grant 2102832. G.V., H.B., J.P., and B.M. acknowledge funding from ERC StG ComplexSwimmers. C.M. acknowledges support through grant RYC-2015-17896 funded by MCIN/AEI/10.13039/501100011033 and “ESF Investing in your future”, grants BFU2017-85693-R and PID2021-125386NB-I00 funded by MCIN/AEI/10.13039/501100011033/ and “ERDF A way of making Europe”.
Competing interest
The authors declare no competing interests.
Method Output level Type of Approach Diffusion parameters Diffusion classes
Hidden Markov method with measurement noise propagation [87, 88] Single trajectory Stats $D$ confined/Brownian
Hidden Markov method and simulations [43] Single trajectory Stats $D$, $\alpha$
Supervised trajectory segmentation algorithm with supported vector machine [89] Single trajectory ML confined/drift/directed
Deep learning and moment scaling spectrum [90] Single trajectory ML + Stats $D$ immobile/Brownian (fast/slow)
Piecewise linear approximation [91] Single trajectory Stats velocity
Detection of velocity and diffusion coefficient[92, 93] Single trajectory Stats $D$, velocity Brownian/directed
Back-propagation neural network [94] Single trajectory ML confined/directed
Divide-and-conquer moment scaling spectrum [95] Single trajectory Stats $D$, $\alpha$, confinement radius immobile/confined/Brownian/directed
Local convex hull [96] Single trajectory Stats $D$,$\alpha$
Recurrence analysis [97] Single trajectory Stats $D$,$\alpha$
Gamma mixture and a hidden Markov method [98] Single trajectory Stats + ML $D$
Superstatistical framework [99] Single trajectory Stats activity, persistence
Probability distribution of square displacements [100] Ensemble Stats $D$, $\alpha$ immobile/confined/Brownian/directed
Probability of transient confinement zones [101] Single trajectory Stats confined/Brownian
Probability of jumps and transient confinement zones [102] Single trajectory Stats confined/Brownian + jumps
Packing coefficient [103] Single trajectory Stats immobile/confined/Brownian
Nonparametric Bayesian inference [104] Ensemble Stats + ML $D$
Displacement distribution and autocorrelation of squared displacements [105] Ensemble Stats $D$
Hierarchical Dirichlet process modeling [106] Single trajectory Stats force
Sliding window of time-averaged MSD [107] Single trajectory Stats $D$,$\alpha$
Random forest [108] Single trajectory Stats confined/Brownian/directed/anomalous
Gyration quantification and a Bayesian statistics [109] Single trajectory Stats $D$ fast/slow
Variational Bayesian treatment of hidden Markov method [110] Ensemble Stats $D$
Bayesian model selection of hidden Markov method [111] Single trajectory Stats $D$ Brownian/directed
Measurement of anomalous diffusion using recurrent neural networks [37] Single trajectory ML $\alpha$
Pointwise diffusion properties with transformers [45] Single trajectory ML $D$ and $\alpha$
Graph-neural network with attention [40] Single trajectory & Ensemble ML $D$, $\alpha$
Recurrent neural network-based autoencoders [112] Single trajectory ML $D$, $\alpha$
Semantic segmentation with convolutional networks [113] Single trajectory ML $\alpha$
Table 1: Methods for the analysis of heterogeneous diffusion. Methods for the analysis of heterogeneous diffusion are classified based on the level of output provided (single-trajectory methods vs. ensemble methods) and on the type of approach (machine learning (ML), classical statistics (Stats), or a combination of the two). We also indicate the diffusion parameters that each method estimates and the diffusion classes that each method uses to classify the trajectories. Only methods accounting for switches of diffusive behavior within the same trajectory are included.
[…]
$\mu$ $\sigma$ $\mu$ $\sigma$
0 SSM 1 1 1 0.01 0.5 0.01 2 -
1 SSM 1 1 0.1 0.01 1.9 0.01 3 -
2 MSM 2 1 $1$ 0.01 1.5 0.01 2 $M=\left(\begin{array}[]{cc}0.99&0.01\\ 0.01&0.99\end{array}\right)$
2 $0.05$ 0.01 0.5 0.01 2
3 MSM 3 1 $1$ $0.01$ 1.5 $0.01$ 2 $M=\left(\begin{array}[]{ccc}0.98&0.01&0.01\\ 0.01&0.98&0.01\\ 0.01&0.01&0.98\end{array}\right)$
2 0.5 0.01 0.5 0.01 2
3 0.01 0.01 0.75 0.01 2
4 QTM 2 1 1 0.01 0.8 0.01 2 $r_{t}=0.6$, $N_{t}=300$, $P_{b}=1$, $P_{u}=0.01$
2 0 0 0 0 0
5 QTM 2 1 1 0.01 1.5 0.01 2 $r_{t}=1.0$, $N_{t}=150$, $P_{b}=1$, $P_{u}=0.05$
2 0 0 0 0 0
6 DIM 2 1 1 0.01 1.2 0.01 2 $r=0.6$, $N=100$, $P_{b}=1$, $P_{u}=0.01$
2 1 0.01 0.8 0.01 2
7 DIM 2 1 1 0.01 1.2 0.01 2 $r=1.0$, $N=80$, $P_{b}=1$, $P_{u}=0.01$
2 3 0.01 0.5 0.01 2
8 TCM 2 1 $1$ 0.01 0.8 0.01 2 $r_{c}=5$, $N_{c}=30$, $T=0.1$
2 $1$ 0.01 0.4 0.01 1
9 TCM 2 1 $1$ 0.01 1 0.01 2 $r_{c}=10$, $N_{c}=30$, $T=0.0$
2 $0.1$ 0.01 1 0.01 1
Table 2: Parameters of the pilot dataset. Example of a dataset composed of 10 experiments reproducing the models of diffusion employed in the $2^{\rm nd}$ Anomalous Diffusion Challenge: single-state model (SSM); multi-state model (MSM); dimerization model (DIM); transient-confinement model (TCM); quenched-trap model (QTM). The diffusion class correspond to $0=$ immobile, $1=$ confined, $2=$ free (unconstrained), $3=$ directed. The parameters specific to each model are: for MSM, the transition matrix $M$; for QTM, the trap radius $r_{\rm t}$, the number of traps $N_{\rm t}$, the probability of trapping $P_{\rm b}$ and untrapping $P_{\rm u}$; for DIM, the interactions radius $r$, the number of particles $N$, the probability of binding $P_{\rm b}$ and unbinding $P_{\rm u}$; for TCM, the compartment radius $r_{\rm c}$, the number of compartments $N_{\rm c}$, and the boundary transmittance $T$. Simulations are provided in generalized units (i.e., pixels and frames) that can be rescaled to meaningful temporal and spatial scales. See A for further details. For all experiments, we simulated $N=100$ particles (Experiment 7 has $N=80$) in a box of size $L=230$ pixel with a FOV size $L_{FOV}=128$ pixel, and a maximum trajectory length of $200$ frames. For Track 1, movies were rendered using a $\mathrm{FWHM}_{\rm PSF}=2.1$ pixel and a $\mathrm{SNR}=7.1$. For Track 2, trajectories were corrupted with Gaussian localization noise with $\sigma_{N}=0.12$ pixel.
Appendix A Simulations of diffusion and interaction models
Trajectories are simulated according to a 2-dimension fractional Brownian motion (FBM) [54]. FBM is a continuous-time Gaussian process $B_{H}(t)$ with stationary increments and a covariance function $E[B_{H}(t)B_{H}(s)]=\frac{1}{2}(|t|^{2H}+|s|^{2H}-|t-s|^{2H})$, where $H$ represents the Hurst exponent and is related to the anomalous diffusion exponent $\alpha$ as $H=\alpha/2$ [54]. The FBM features three regimes: one where the increments are positively correlated ($11/2<H<$ 1, i.e., $21<\alpha<$ 2, superdiffusive); one where the increments are negatively correlated ($20<H<1/$ 2, i.e., $10<\alpha<$ 1, subdiffusive); and one where the increments are uncorrelated ($H=1/2$, i.e., $\alpha=1$, diffusive Brownian motion).

### v3

@@ 4.6 Metrics for challenge ranking
For the Ensemble Task, the metrics involved in the calculation were the F1-score of the model and the MAE of the distributions of $K$ and $\alpha$. For the Single-trajectory Task, we used the JSC and the RMSE of CPs, the MSLE of $K$, and the MAE of $\alpha$.
4.7 Overview of the Challenge results
The Challenge dataset was comprehensively designed to test the submitted methods under distinct scenarios, using ad hoc metrics to evaluate their specific capabilities. For ranking, we employed composite metrics that aggregate the scores from different experiments and subtasks. The results are summarized in Fig.˜3. Here, we present an overview of the Challenge results, highlighting the general trends observed. The complete rankings are provided in Supp. Fig. 2.
In the Single-trajectory Task (Fig.˜3a), one method based on UNet3+ [86, 87] (team I), clearly outperformed the others, whereas the Ensemble Task (Fig.˜3b) showed a more balanced competition. From the MRR breakdown, we observed that the top team in the Single-trajectory Task performed consistently well across all metrics. In contrast, for the Ensemble Task, the top teams improved their final ranks by specializing in one of the two subtasks.
Figure 3: Challenge rankings. a, b Mean reciprocal rank (MRR) of all methods participating in the Single-trajectory Task (a) and Ensemble Task (b) in the Trajectory Track. The colors represent the relative contributions of the metrics of each subtask to the overall MRR. c–e Correlation between subtask metrics associated with changepoint (CP) detection (c), the prediction of segment properties (d) in the Single-trajectory Task, and the prediction of diffusive properties in the Ensemble Task (e). Empty circles and filled squares represent the metrics obtained for the Video Track and the Trajectory Track, respectively. Dashed lines join results obtained by the same team in the two tracks. The darker background color indicates the area of the plot corresponding to better performances. Source data are provided as a Source Data file.
We also show the correlation between pairs of metrics associated with CP detection (Fig.˜3c) and the prediction of diffusive properties (Fig.˜3d,e). The predictions for the Video Track (represented by filled squares) are also included alongside those of the Trajectory Track (represented by empty circles). Across methods, enhanced CP detection, reflected by higher JSC and lower RMSE, yields a tight correlation between these metrics (Fig.˜3c). A similar but weaker trend appears for $K$ and $\alpha$ errors (Fig.˜3d,e), because their estimation often relies on distinct algorithms, decoupling improvements in one from the other.
In the plots, the dashed lines connect the predictions of teams participating in both tracks. All teams in the Video Track (teams E and Q for the Single-trajectory Task, teams E and F for the Ensemble Task), except for team K, improved their predictions in the Trajectory Track compared to the Video Track. Notably, all four teams first extracted the trajectories using a previously established tracking method [88, 89, 90, 5, 40, 91] and then performed the ensuing analysis using the same method developed for the Trajectory Track. While this highlights the influence of error associated with the tracking process [51], none of the methods explored the possibility of obtaining results directly from the video, which was one of the exploratory goals of this competition.
Finally, Fig.˜4 shows the score obtained for subtask metrics by all teams for each experiment (filled symbols). The consistently lower performance of the Video Track compared to the Trajectory Track lends support to the third rationale: it suggests that challenges in accurately extracting trajectories from experimental videos represent a more significant bottleneck than the downstream analysis of pre-extracted tracks.
Figure 4: Overview of the results. a–e Scores obtained for each subtask metric by all teams for each experiment of the Single-trajectory Task (filled symbols). Squares and circles represent the metrics obtained for the Video Track and the Trajectory Track, respectively. f,g Scores obtained for each subtask metric by all teams for each experiment of the Ensemble Task (filled symbols). Squares and circles represent the metrics obtained for the Video Track and the Trajectory Track, respectively. In all panels, open symbols represent average scores. The vertical axes are arranged so that the best performance is always shown at the top. Horizontal square brackets indicate groups of experiments that are discussed comparatively. Source data are provided as a Source Data file.
These plots provide further insight into which experimental conditions were more challenging for each subtask. For example, CP detection in EXP 1 (MSM with 3 states) was particularly difficult, as indicated by the low JSC in Fig.˜4a. As shown in Fig.˜4e, classification of the type of diffusion for EXP 4 (TCM) was more challenging than EXP 3 (QTM), despite having similar parameters for the unrestrained motion. For the Ensemble Task, we observe poorer predictions for $K$ in EXP 8 (SSM, Fig.˜4f) and for $\alpha$ in EXP 9 (QTM, Fig.˜4g). In the following, we will comparatively discuss results obtained for groups of experiments aimed at detecting specific method capabilities. For most of these analyses, we will mainly consider the methods of the top 5 teams in each Track and Task.
4.8 CP detection and segment diffusion properties
A main aspect of the Challenge was the evaluation of CP detection capability and the ensuing assessment of diffusion properties for the identified segments. In particular, we tested the methods’ ability to distinguish true anomalous diffusion from subdiffusive behavior that emerges solely from physical constraints, directly addressing the second rationale. These insights were provided by the Single-trajectory Task.
Figure 5: CP detection and segment diffusion properties. a, b Scatter plots of the metrics associated with segment diffusion properties in the Single-trajectory Task of the Trajectory Track as a function of the false positive rate calculated for EXP 8 composed of trajectories without CPs. Results obtained by all participants are shown. Dashed lines correspond to the scores obtained by using the fit of the mean-squared displacement, as a benchmark. The darker background color indicates the area of the plot corresponding to better performances. c–e Dependence of the JSC, MSLE, and MAE on the number of changepoints for the Single-trajectory Task of the Trajectory Track. Only the results of the top 5 teams are shown. The color code represents their position in the ranking (blue is the highest, red the lowest). Source data are provided as a Source Data file.
As shown in Fig.˜4a–e, the methods generally performed well when tested on time-varying processes. We sought to characterize the false positive rate of the methods by evaluating their behavior over the trajectories of EXP 8 having no CPs (Fig.˜5a,b). EXP 8 also served to assess the methods’ ability to estimate parameters $K$ and $\alpha$ independently of errors induced by incorrect segmentations. Submitted predictions were benchmarked with the estimations of $K$ and $\alpha$ obtained by linear and logarithmic fits of the MSD, respectively (dashed lines). Most methods predicted very few CPs for these trajectories, producing a low false positive rate and outperformed the MSD fit for both $K$ and $\alpha$ (Fig.˜5a,b).
A relevant aspect associated with CP detection accuracy is its dependence on the number of CPs per trajectory, shown in Fig.˜5c–e, which is inversely related to the average segment duration. As expected, the JSC shows worse performance as the number of CPs increases (Fig.˜5c). Regarding the diffusion parameter estimation, we observe that the methods allow a robust estimation of $K$ independently of the number of CPs (Fig.˜5d), whereas for $\alpha$ we observe a drop in performance as the number of CPs increases (Fig.˜5e). This confirms the difficulty of estimating $\alpha$ from short segments, due to its asymptotic nature, already observed in the 1st AnDi Challenge [42].
4.9 Classification of types of diffusion
One of the goals of this competition was to assess the methods’ ability to classify different diffusion types and distinguish among distinct physical models. Results for all experiments of the Video and Trajectory Tracks are shown in Supp. Fig. 3 and Supp. Fig. 4, respectively. The results of the two tracks were qualitatively similar but the Video Track had overall lower scores since all teams except team Q missed the immobile state (Supp. Fig. 3). To summarize the methods’ ability to assign segments to diffusion types, in Fig.˜6 we show the distribution of each diffusive state compared to the ground truth (horizontal segments) for representative experiments of the Trajectory Track. In Fig.˜6a we exemplarily show the results obtained for EXP 9, a QTM with an unconstrained state having a narrow distribution of $K$ but with $\alpha$ values that could produce either superdiffusive or directed motion. In this case, only the top method (team I, light blue) was able to produce a reliable classification of the diffusion type of the segments. The difficulty in inferring the correct type of mechanism producing interaction underscores the challenges in accurately analyzing this kind of data, which can have significant implications for the biological interpretation of the results. Although perfect classification of diffusive states remains challenging, the algorithms nonetheless provide precise estimates of critical biophysical parameters, namely, the average dwell times in both trapped and unconstrained states (inset of Fig.˜6a). The measure of these parameters is essential for quantifying binding kinetics, confinement lifetimes, and transition rates that directly inform biological interpretation.
Figure 6: Classification of types of diffusion. Only the results of the top 5 teams are shown, the color code represents their position in the ranking (blue is the highest, red the lowest). a–d Predictions for the frequency of time spent with a given diffusion type for EXP 9 (a), EXP 3 (b), EXP 4 (c) and EXP 5 (d) for the Single-trajectory Task of the Trajectory Track. Horizontal segments represent the ground truth. Inset of panel a: predicted average residence time for the trapped and non-trapped states. The non-trapped state includes segments corresponding to both unconstrained (free/anomalous) diffusion and directed motion. Source data are provided as a Source Data file.
The second rationale for the Challenge was to probe the methods’ ability to disentangle genuine anomalous diffusion from subdiffusive behaviors arising purely from motion constraints. To test the methods in challenging conditions, we designed a group of experiments (EXP 3, EXP 4, and EXP 5) with different underlying models but with diffusive parameters that produce similar trajectories. The three experiments share an unconstrained state with normal diffusion and $K\approx 1$: EXP 3 is simulated as QTM, whereas EXP 4 is from a TCM with a small confinement radius and $\alpha\approx 0.2$, and EXP 5 is DIM with a dimeric state with $\alpha\approx 0.2$. Other parameters were set to obtain similar residence times in the different states. Fig.˜6b–d highlights the performance of the top five methods across EXP 3–5. Teams I, C, and R each correctly classify over 95% of segments, closely matching the true distribution of diffusive states. Team E tends to over-label segments as diffusive, while Team O occasionally confuses confined segments for diffusive ones and vice versa. Team R, despite its high overall accuracy, also makes occasional misclassifications of diffusive segments as immobile or confined. Importantly, for EXP 4 (small-radius confinement) and EXP 5 (dimerization-induced subdiffusion), misclassification as immobile is negligible for Teams I, C, and R. Detecting confinement in EXP 4 is particularly challenging since short dwell times in confined areas yield few boundary reflections, inducing confusion with unconstrained anti-persistent subdiffusion of EXP 5. The ability of Teams I, C, and R to resolve these subtle cases underscores the high sensitivity and robustness of their methods.
4.10 Using physical models to enhance method performance
The information contained in an individual trajectory is typically sufficient to estimate CPs and diffusive properties. However, for some physical models, the knowledge of the model itself offers additional information that could be used to improve further CP detection and parameter estimation. This is the case for QTM and TCM, where changes in diffusion correspond to spatial constraints. For DIM, diffusion changes are associated with particle proximity; in addition, since particles in a dimer co-diffuse, one could in principle use twice as much information to estimate $K$ and $\alpha$, although in typical experimental conditions it may be very challenging to track two co-diffusing particles.
Along these lines, for the Single-Trajectory Task, the lowest JSC values were obtained for EXP 1 and EXP 7 (circles in Fig.˜4a). Both experiments correspond to simulations of MSM, a model where the diffusion changes are produced in a purely time-dependent fashion and the dataset itself does not provide additional hints to determine them. This suggests that the methods can directly or indirectly take advantage of the presence of a physical event (e.g., trapping, confinement, or dimerization) to enhance CP detection accuracy. To assess this effect quantitatively, we used EXP 5 and EXP 6, which correspond to different physical models (DIM and MSM, respectively) generated with an identical set of diffusive parameters. To quantify model-based gains, we computed the relative improvement
$\Delta m(\%)=\frac{m_{\mathrm{DIM}}-m_{\mathrm{MSM}}}{m_{\mathrm{MSM}}}\times 100\%$ (14)
for each subtask metric (JSC, MSLE, and MAE). Fig.˜7 reports these improvements for all methods, with the overall average shown as a dashed line.
Figure 7: Effect of the physical model. Relative improvement for all participants when predicting trajectory properties from EXP 5 (DIM) with respect to EXP 6 (MSM) for the Single-trajectory Task of the Trajectory Track. Each panel reports the metrics associated with a different subtask: JSC (a), MSLE (b), and MAE (c). Source data are provided as a Source Data file.
Surprisingly, while most of the methods showed improved performance for the CP prediction in DIM (Fig.˜7a), there were minor differences in the prediction of diffusive properties (Fig.˜7b,c). We believe this is because the methods predict each trajectory’s properties without considering it in the ensemble of the FOV or of the experiment, an observation that may improve the next generation of methods.
4.11 Ensemble predictions
The Ensemble Task was designed to test whether the methods could take advantage of the increased statistics obtained from common parameters shared by all trajectories within the same experiment to better identify the type of motion and estimate its parameters. As discussed earlier, several approaches of this type have been devised and used in the past to extract biophysical information from single-particle tracking data (Supp. Tab. 1). However, no pure ensemble-level method, i.e., one that disregards the individual trajectory identity, was employed for the Challenge. Instead, all teams that provided submissions for the Ensemble Task used predictions obtained at the single-trajectory level, which were then pooled together to estimate the moments of the distributions of the diffusive parameters. Results for all experiments of the Video and Trajectory Tracks are shown in Supp. Figs. 5-8. The resulting distributions are summarized in Fig.˜8 for 4 exemplary experiments (EXP 4, EXP 7, EXP 8, and EXP 9) of the Trajectory Track. The pooling operation was performed using two general approaches: teams either applied a Gaussian mixture model (GMM) or a clustering algorithm on the predicted segments to extract subpopulation parameters, with four of the top 5 teams opting for the former approach (teams E, I, M, and O). Interestingly, as it can be inferred from Fig.˜3a,b, the scores obtained by the teams participating in both tasks showed a low correlation. Therefore, accurate predictions at the single-trajectory level do not necessarily translate into reliable ensemble-level predictions, pointing to a critical role of the clustering approach.
Figure 8: Ensemble Task predictions for the Trajectory Track. Panels a–d show the predicted distributions of the diffusion coefficient $K$ and panels e–h the anomalous exponent $\alpha$ for EXP 4, EXP 7, EXP 8, and EXP 9, respectively. Distributions were computed from the estimated means and variances (see Scoring and evaluation – Ensemble Task). Only the results of the top five teams are displayed, with color indicating rank (blue is the highest, red the lowest). Black curves denote the ground-truth distributions. Source data are provided as a Source Data file.
Fig.˜8a,b show an experiment where all teams provided consistent and reasonable predictions. This is particularly evident for the $K$ distribution in EXP 7 and EXP 8 (Fig.˜8b,c). Since the methods rely on estimates of $K$ per segment and then apply GMM or k-means, they generally tend to over-fragment wide $K$ ranges, misrepresenting the overall distribution. The corresponding predictions for the distributions of $\alpha$ for these experiments are shown in Fig.˜8f,g. For EXP 8, characterized by the absence of CPs and nearly flat distributions of $K$ and $\alpha$, most methods successfully captured the broad distribution of $\alpha$ (Fig.˜8g). However, their predictions for $K$ (Fig.˜8c) were often biased toward different ranges within the allowed support. In contrast, EXP 9 presented a population of short dwell times in the trapped state. Most methods successfully detected the occurrence of these events, as reflected in the $K$ distribution (Fig.˜8d) but, with the exception of team I, failed to associate these events with the correct $\alpha=0$ Fig.˜8h.
We further point out that optimizing methods to provide high scores for the metrics of the competition did not always translate into more meaningful insights about the underlying physical processes. For instance, teams M, H, and O showed significant biases across all experiments when predicting the $K$ distribution but still achieved high rankings according to the metric in Eq.˜3 (Supp. Fig. 6). Moreover, accurately predicting the number of true states did not provide a clear advantage with this metric, as most top teams overestimated the number of states but carefully adjusted their relative weights to minimize differences with the ground-truth distribution.
4.12 Results Summary and Take-home Messages
Robust changepoint detection:
Top single-trajectory methods (e.g., based on UNet3+ [86]) consistently achieve over 95% accuracy in identifying segment boundaries, with only minor false-positive rates across all scenarios.
Distinguishing confinement, immobilization, and anomalous diffusion:
Leading algorithms accurately classify segments arising from geometric constraints or anomalous dynamics. Only very short segments and exponents close to zero remain challenging, indicating minimal crosstalk between distinct diffusion mechanisms.
Trajectory extraction is a bottleneck:
Video-Track performance lags the Trajectory Track by 10–30%, highlighting that linking and localization errors—not downstream analysis—drive most of the accuracy loss.
Parameter estimation benefits from physical priors:
Incorporating known physical models may yield significant gains in changepoint detection, but separate estimation pipelines for $K$ and $\alpha$ result in only modest improvements in parameter accuracy.
Dedicated ensemble approaches are needed:
Ensemble Task submissions rely on GMM or k-means clustering of per-trajectory outputs, which fragments broad parameter distributions (e.g., EXP 7–8). Ensemble approaches, either bypassing single-trajectory clustering or using more sophisticated grouping techniques, hold potential for uncovering population-scale insights.
5 Discussion
The 2nd AnDi Challenge provided a platform for advancing methods to characterize diffusion trajectories, with a special focus on those exhibiting transitions between distinct diffusive regimes. Through this Challenge, participants developed approaches that, when applied to standardized benchmarks, demonstrate robust capabilities in analyzing processes akin to those found in complex biophysical environments.
The high participation from teams spanning different fields vividly demonstrated the first rationale for the Challenge: the urgent need for standardized, rigorously evaluated methods to analyze dynamic changes in particle motion.
The Challenge highlighted several key insights. The methods for changepoint analysis have reached a good level of maturity. Participants demonstrated strong capabilities in detecting changepoints, which is crucial for understanding transitions between different diffusive regimes. However, the characterization of the resulting segments can still be improved. Accurate estimation of diffusion parameters within these segments remains challenging, particularly for short segments where the asymptotic nature of certain parameters, such as the anomalous diffusion exponent $\alpha$, complicates analysis. Sequence-to-sequence machine learning methods, mostly based on architectures combining convolutional [92] and transformer [93] layers, have shown great flexibility and effectiveness. The top-performing methods often utilized these architectures, highlighting their potential for further advancements in the field. Notably, the methods did not take into account information coming from common parameters shared among trajectories or the underlying physical processes. Incorporating this knowledge could enhance the accuracy and robustness of the analyses.
Nevertheless, significant challenges remain, and we hope the Challenge will help pave the way toward their resolution. In particular, we highlight two promising new avenues, which we believe may have a great impact on our understanding of the physics underlying biophysical processes.
First, the precision with which we can extract diffusion parameters remains fundamentally limited by current tracking algorithms, directly highlighting the third rationale for the Challenge. Notably, all participants in the Video Track relied on existing tracking techniques, subsequently applying the methods developed for the Trajectory Track to analyze the resulting trajectories. Despite the rapid advances in deep learning, none of the participants have yet leveraged these cutting-edge technologies to directly extract diffusive properties from video data. This missed opportunity could be attributed to several factors: the analysis technology may not yet be fully mature, the training processes might be too lengthy and complex, or the computational resources and time required could be prohibitively high. We foresee that as these bottlenecks are addressed, a new generation of methods will emerge, capable of bypassing the tracking step altogether and setting new standards of accuracy.
Second, in the Ensemble Task, all participants relied on post-processing of single-trajectory outputs. Features were first extracted from individual trajectories and then a separate step was used to infer the parameters of the diffusive populations. No team developed dedicated ensemble-level algorithms or used established ensemble frameworks.
Although this single-trajectory-based approach produced high Challenge rankings, it offered limited biophysical insight due to the proliferation of predicted states and the instability of each mode’s variance. Minimizing the Wasserstein-1 ($W_{1}$) distance aligns predicted and ground-truth distributions, but $W_{1}$ offers no penalty for over-splitting into numerous states or for unstable variance estimates, nor does it encourage physically interpretable solutions (e.g., filtering overlapping modes or very low-population segments). This warns us that outputs should not be blindly trusted when applied to real experiments. Care should always be taken not to overfit the data with too many states that cannot be assigned to a biophysical process. Whenever an analysis yields a large number of states, their identities should be validated through control experiments. In practice, a priori biological knowledge often narrows the expected state count, providing essential context for interpreting algorithmic results.
Looking ahead, methods capable of inferring population distributions directly from the raw ensemble of trajectories, thereby bypassing single-trajectory feature extraction and clustering, may deliver deeper physical insights. Moreover, approaches that treat the full set of trajectories contextually, rather than in isolation, are likely to enhance both performance and interpretability.
To encourage further development of methods addressing these issues, as well as those aligned with approaches used throughout the challenge, we have made the labeled dataset discussed in this work publicly available on Zenodo [94]. This resource allows researchers to benchmark new methods in a standardized manner, while also providing the experimental biophysics community with a tool to better identify the methods best suited to their specific experimental scenarios.
6 Methods
6.1 Simulations of diffusion and interaction models
Trajectories are simulated according to a 2-dimensional fractional Brownian motion (FBM) [54]. FBM is a continuous-time Gaussian process $B_{H}(t)$ with stationary increments and a covariance function $E[B_{H}(t)B_{H}(s)]=\frac{1}{2}(|t|^{2H}+|s|^{2H}-|t-s|^{2H})$, where $H$ represents the Hurst exponent and is related to the anomalous diffusion exponent $\alpha$ as $H=\alpha/2$ [54]. FBM features three regimes: one in which the increments are positively correlated ($1/2<H<1$, i.e., $1<\alpha<2$, superdiffusive); one in which the increments are negatively correlated ($0<H<1/2$, i.e., $0<\alpha<1$, subdiffusive); and one in which the increments are uncorrelated ($H=1/2$, i.e., $\alpha=1$, diffusive Brownian motion).